OpenStudy (anonymous):

While traveling south on an expressway, a car traveling 60 mph (miles per hour) slows down to 30 mph in 12 minutes due to traffic conditions. Calculate the acceleration

OpenStudy (anonymous):

I have a=-2.5 mi/hsquared S

OpenStudy (anonymous):

Is my answer correct?

OpenStudy (anonymous):

simple formula v = u - at put v = 30 u = 60 t = 12*60

OpenStudy (anonymous):

30=60(1-12a)

OpenStudy (anonymous):

12a =.5

OpenStudy (anonymous):

a = 1/24 =0.0416 in the negative x direction

OpenStudy (anonymous):

that's the answer?

OpenStudy (anonymous):

is my answer wrong ??

OpenStudy (anonymous):

all the possible answers are whole numbers

OpenStudy (anonymous):

then probably the answers are wrong because u and i both know that v = u-at is the formula unless any different condition is there in the question . so did u type the question correctly

OpenStudy (anonymous):

no im i thought i saw periods sorry. the answers are -150 -6.0 -450 -2.5

OpenStudy (anonymous):

oh ho !! got it . your answers are given in hours . i considered speed in miles/hour but time i made it in seconds !!

OpenStudy (anonymous):

correct ans 30 = 60 - 12a a = 30/12 = 2.5 in negative x direction

OpenStudy (anonymous):

no im i thought i saw periods sorry. the answers are -150 -6.0 -450 -2.5

OpenStudy (anonymous):

uff !! i am again wrong correct ans 30 = 60 - 12/60a a = 30* 60/12 = 150 mi/h^2

OpenStudy (anonymous):

this time i really got the ans .. after dividing 12/60 i just converted 12min in hours

OpenStudy (anonymous):

which one would it be out of those 4?

OpenStudy (anonymous):

-150

OpenStudy (anonymous):

thanks